# A quantum computing concept for 1-D elastic wave simulation

### v1

@@
A quantum computing concept for 1-D elastic wave simulation
[…]
${}^{1}$ Department of Earth and Planetary Sciences
@@
   Switzerland. Email: cyrill.boesch@erdw.ethz.ch
${}^{\dagger}$ These authors contributed equally
@@
Quantum computing has attracted considerable attention in recent years because it promises speed-ups that conventional supercomputers cannot offer, at least for some applications. Though existing quantum computers are, in most cases, still too small to solve significant problems, their future impact on domain sciences is already being explored now. Within this context, we present a quantum computing concept for 1-D elastic wave propagation in heterogeneous media with two components: a theoretical formulation and an implementation on a real quantum computer. The method rests on a finite-difference approximation, followed by a sparsity-preserving transformation of the discrete elastic wave equation to a Schrödinger equation, which can be simulated directly on a gate-based quantum computer. An implementation on an error-free quantum simulator verifies our approach and forms the basis of numerical experiments with small problems on the real quantum computer IBM Brisbane. The latter produce simulation results that qualitatively agree with the error-free version but are contaminated by quantum decoherence and noise effects. Complementing the discrete transformation to the Schrödinger equation by a continuous version allows the replacement of finite differences by other spatial discretisation schemes, such as the spectral-element method. Anticipating the emergence of error-corrected quantum chips, an analogy between our method and analyses of coupled mass-spring systems suggests that our quantum computing approach may lead to wave field simulations that run exponentially faster than simulations on classical computers.
@@ 1 Introduction
QCs harness quantum phenomena such as entanglement and superposition for problem-solving algorithms, thereby promising up to exponential reductions in both runtime and memory usage compared to some classical algorithms (e.g., Nielsen and Chuang, 2010; Benenti et al., 2004). As a result, quantum computation could make it feasible to tackle classes and sizes of problems considered intractable on traditional computing hardware (e.g., Shor, 1994; Harrow et al., 2009; Montanaro, 2016). Recent studies have started to explore quantum computing to address geophysical problems by proposing quantum simulation algorithms for solving the acoustic wave and Maxwell’s equations (e.g., Moradi et al., 2018, 2019; Suau et al., 2021; Jin et al., 2022, 2023). In a broad sense, seismic imaging and related uncertainty quantification problems continue to be primarily limited by computational resources. It is, therefore, timely to assess the potential of quantum computing for the simulation of seismic wave propagation in heterogeneous media, despite the immaturity of the existing QCs.
This study is intended to be an early step in this direction. We show how elastic wave propagation in heterogeneous media can be simulated using a QC and perform initial experiments on actual QCs. More specifically, we extend numerical wave propagation via Hamiltonian simulation (Costa et al., 2019; Suau et al., 2021) to heterogeneous media. For this, we employ an analytical Cholesky decomposition, which facilitates the transformation of the elastic wave equation with heterogeneous medium properties to a Schrödinger equation. While we employ a finite-difference approximation of the elastic wave equation, we also discuss the general continuous case and how other numerical methods can be applied. We proceed with the implementation of our method on the QC IBM Brisbane, in order to solve a small-scale problem and explain current challenges. These include systematic errors caused by noise and decoherence rates, which are characteristic for current NISQ QCs. Anticipating future developments, we finally delve into the potential scalability of an optimal Hamiltonian simulation algorithm on future quantum hardware equipped with error correction. By mapping our problem to a mass-spring model analysed in Babbush et al. (2023), we establish that the algorithm is capable of providing exponential speedup compared to computations on classical computers.
@@ 2 A brief summary of quantum computing
To set the stage, we begin with a condensed review of the tensor product formalism and the quantum circuit model that should be accessible to a geophysical audience. This framework is inherently dictated by the postulates of quantum mechanics and forms the fundamental programming model of quantum computing. A more general formalism of density matrices, as well as more in-depth introductions, can be found in the literature (Nielsen and Chuang (2010); Watrous (2018); Adedoyin et al. (2022)).
@@ 2.1 Quantum particles and state space
Consider a single quantum particle characterised by its quantised properties, such as the atomic energy level. This particle can occupy $N$ different orthogonal basis states, which span the state space of the particle. To the basis states, we assign labels $\Sigma=\{0,\dots,N-1\}$, collectively referred to as the label set of size $|\Sigma|=N$. A particle with $N$ basis states is called an $N$-level particle. The current state of the observable is represented by the state vector |$\ket{\psi}=[\alpha_{0},\dots,\alpha_{N-1}]^{T}\in\mathbb{C}^{N}$, where $\alpha_{k}\in\mathbb{C}$ is the probability amplitude corresponding to the $k$-th basis state |$\ket{k}$, written in the bra-ket or Dirac notation that is ubiquitous in quantum mechanics. It encapsulates all information about the individual particles and their number and concisely represents exponentially large state vectors. The ket |$\ket{a}$ is a column vector, whereas the bra $\bra{b}$ | is a row vector with complex conjugate transpose $\bra{b}^{\dagger}=\ket{b}$. Bra and ket form the bra-ket $\braket{b}{a}$, which is the scalar product |$\ket{b}^{\dagger}\cdot\ket{a}$. The main difference, compared to the common vector notation in linear algebra, is that the content of the bra/ket does not represent individual vector entries but some convenient collective label. The state vector |$\ket{\psi}$ can be expressed in terms of the basis states |$\ket{k}$ with $k=0,...,N-1$ as the linear combination |$\ket{\psi}=[\alpha_{0},\dots,\alpha_{N-1}]^{T}=\alpha_{0}\ket{0}+\alpha_{0}\ket{1}+\dots+\alpha_{N-1}\ket{N-1}\in\mathbb{C}^{N}$, which is called a superposition state.
@@
where |$\ket{\psi_{j}}\in\mathbb{C}^{N_{j}}$ is the state of the $j^{\text{th}}$ particle, and $N_{j}$ is the number of states that it can occupy. We reuse the symbols $N$ and $\psi$, because a single-particle system is a special case of a multi-particle system. The symbol $\otimes$ denotes the Kronecker or tensor product. For two arbitrary states $\mathbf{u}$ and $\mathbf{v}$, it is defined as
@@
Any quantum state must be normalised, i.e., satisfy the normalisation condition
@@
Combining normalised states with the $\otimes$ product, the resulting composite state is normalised as well. If $\alpha_{k}\neq 0$ for at least two different indices $k$, then the state |$\ket{\psi}$ simultaneously holds multiple potential outcomes $\Omega=\{\ket{k}:\alpha_{k}\neq 0\}$ and is called a superposition state. In contrast to classical systems, a measurement of the quantum system causes the state |$\ket{\psi}$ to collapse from the superposition into one of the possible basis states |$\ket{k}\in\Omega$ randomly with probability $P(k)=|\alpha_{k}|^{2}$, which explains the term probability amplitude used for the coefficients $\alpha_{k}$. The type of measurement just introduced is a full-system measurement in the standard basis, which is the simplest case; we do not deal with more general cases, which can be found in the literature at the top of this section. Measurements allow us to read the label of the collapsed state in the form of classical bits, forming the only interface between quantum and classical information.
@@ 2.3 Hamiltonian simulation
with the initial state |$\ket{\psi(0)}$. The time evolution operator $\mathbf{U}(t)=\exp(-i\mathbf{H}t)$ is unitary. Our approach to elastic wave propagation on a QC rests on the quantum or Hamiltonian simulation (HS) approach (e.g., Georgescu et al., 2014; Nielsen and Chuang, 2010), which amounts to the direct evaluation of (9). HS was originally introduced to model multi-particle quantum dynamic systems using QCs. Due to the aforementioned exponential space growth, quantum dynamics modelling becomes prohibitively complex on classical computers. In contrast, by being quantum systems themselves, QCs naturally feature the same state space expansion, potentially making multi-particle simulations possible.
For initialisation, HS requires the application of a unitary operation $\mathbf{V}$ that sets the desired initial state, i.e. |$\ket{\psi(0)}=\mathbf{V}\ket{00..00}$. Subsequently, the unitary time evolution operator $\mathbf{U}(t)$, constructed using basic gates, is applied to |$\ket{\psi(0)}$,
@@
The main challenge of HS is that the matrix exponential $\mathbf{U}(t)=\exp(-i\mathbf{H}t)$ is notoriously difficult to evaluate explicitly. Hence, the application of $\mathbf{U}(t)$ is replaced by some unitary $\mathbf{U}_{\textrm{A}}(t)$ that evolves the initial state to the final one with the desired error $\varepsilon$, $||\mathbf{U}(t)\ket{\psi(0)}-\mathbf{U}_{\textrm{A}}(t)\ket{\psi(0)}||<\varepsilon$, and can be implemented efficiently using basic quantum gates, with circuit depth growing along with $1/\varepsilon$.
@@
Other approaches can be found in the literature, with efficiency varying based on properties of the given $\mathbf{H}$. These include methods based on the Trotter-Suzuki Decomposition (e.g., Trotter, 1959; Suzuki, 1976, 1991; Hatano and Suzuki, 2005; Dhand and Sanders, 2014; Yi and Crosson, 2022) and the recent qubitisation approach (Low and Chuang, 2019), discussed more in detail in appendix B.
It has been shown that HS can achieve up to exponential speedup compared to analogous classical algorithms (e.g., Berry et al., 2007; Childs et al., 2018; Daley et al., 2022; Babbush et al., 2023). We will discuss this matter as well as the question of utilising the final state in section 5.
@@ 3 Elastic waveform propagation using HS
with time $t$, the wave field $u(x,t)$, density $\rho(x)$ and elastic modulus $\mu(x)$. Furthermore, the wave field is subject to the initial conditions $u_{0}=u(0,x)$ and $v_{0}=\partial_{t}u(0,x)$. We introduce a first-order finite-difference discretisation in space with $2^{n-1}$ grid points, each with position and velocity, and a grid spacing of $\Delta x=W/(2^{n-1}-1)$, thereby approximating the continuous wave field $u(x,t)$ by the discrete wave field vector $\mathbf{u}(t)=[u(x_{0},t),\dots,u(x_{2^{n-1}-1},t)]^{T}$. In section 5.3, we discuss the potential use of other discretisation methods. We further define the positive definite mass matrix $\mathbf{M}$ that serves as a discrete representation of the density distribution,
@@
This implementation naturally results in a Dirichlet and a Neumann boundary condition on the right and left, respectively, assuming two ghost nodes. In the interest of a smooth derivation, we defer the case of arbitrary boundary conditions to appendix C. With these definitions, the space-discretised wave equation reads
@@
To solve (16) on a QC, it needs to be transformed into the Schrödinger equation (8). For this, we first apply a transformation that utilises the positive definite, diagonal square root matrix $\mathbf{M}^{1/2}$, which produces the mass-transformed wave field $\tilde{\mathbf{u}}=\mathbf{M}^{1/2}\mathbf{u}$ and the mass-transformed stiffness matrix $\tilde{\mathbf{K}}=\mathbf{M}^{-1/2}\mathbf{K}\mathbf{M}^{-1/2}$. The latter inherits the property of being symmetric negative-definite from $\mathbf{K}$. We proceed by transforming the second-order system into a first-order system
@@
where $\tilde{\mathbf{v}}=\partial_{t}\tilde{\mathbf{u}}=\mathbf{M}^{1/2}\partial_{t}\mathbf{u}\in\mathbb{R}^{2^{n-1}}$ is the mass-transformed velocity vector and $\mathbf{Q}$ the impedance matrix. While (17) is already closer to the Schrödinger equation than (16), it remains to symmetrise $\mathbf{Q}$ and obtain the imaginary unit $i$ on the left-hand side. For this, we multiply (17) by an arbitrary invertible transformation $\mathbf{T}\in\mathbb{R}^{2^{n}\times 2^{n}}$, which results in
@@
which are typically sparse in seismic applications thanks to the spatial localisation of sources. Furthermore, owing to the localised interactions of the discretised degrees of freedom, both the mass and stiffness matrices are sparse, which implies that $\mathbf{Q}$ is sparse, too. Eq. (19) can, in principle, be solved by a QC, and it is isomorphic to the discretised elastic wave equation (16) because $\mathbf{T}$ is invertible. The actual elastic wave field at all times can, therefore, be reconstructed via

### v2

@@
A quantum computing concept for 1-D elastic wave simulation with exponential speedup
Malte Schade,1,† Cyrill Bösch,1,† Václav Hapla,1 Andreas Fichtner1
1 Department of Earth and Planetary Sciences
@@
   Switzerland. Email: cyrillboesch@gmail.com
These authors contributed equally
@@
Quantum computing has attracted considerable attention in recent years because it promises speedups that conventional supercomputers cannot offer, at least for some applications. Though existing quantum computers are, in most cases, still too small to solve significant problems, their future impact on domain sciences is already being explored now. Within this context, we present a quantum computing concept for 1-D elastic wave propagation in heterogeneous media with two components: a theoretical formulation and an implementation on a real quantum computer. The method rests on a finite-difference approximation, followed by a sparsity-preserving transformation of the discrete elastic wave equation to a Schrödinger equation, which can be simulated directly on a gate-based quantum computer. An implementation on an error-free quantum simulator verifies our approach and forms the basis of numerical experiments with small problems on the real quantum computer IBM Brisbane. The latter produce simulation results that qualitatively agree with the error-free version but are contaminated by quantum decoherence and noise effects. Complementing the discrete transformation to the Schrödinger equation by a continuous version allows the replacement of finite differences by other spatial discretization schemes, such as the spectral-element method. Anticipating the emergence of error-corrected quantum chips, we analyze the computational complexity of the best quantum simulation algorithms for future QCs. This analysis suggests that our quantum computing approach may lead to wave field simulations that run exponentially faster than simulations on classical computers.
@@ 1 Introduction
QCs harness quantum phenomena such as entanglement and superposition for problem-solving algorithms, thereby promising up to exponential reductions in both runtime and memory complexity compared to some classical algorithms (e.g., Nielsen and Chuang, 2010; Benenti et al., 2004). As a result, quantum computation could make it feasible to tackle classes and sizes of problems considered intractable on traditional computing hardware (e.g., Shor, 1994; Harrow et al., 2009; Montanaro, 2016). As of today seismic imaging and related uncertainty quantification problems continue to be primarily limited by computational resources. It is therefore natural and timely to ask whether the potential of QCs can be harnessed for such inverse problems (Moradi et al. (2018, 2019)). Recent studies have used adiabatic quantum computers (O’Malley (2018); Golden and O’Malley (2021); Souza et al. (2022); Dukalski et al. (2023)) and variational quantum algorithms (Trahan et al. (2023)) for solving geoscience-related problems (see Section 5.5 for more details). In all studies, the quantum resources are used to solve linearised inverse problems. Here, we intend to take a first step toward non-linear Full Wave Form Inversion (FWI) (Virieux and Operto (2009); Fichtner (2010); Liu and Gu (2012)) by tackling the forward problem of solving the elastic wave equation, using a different algorithmic approach, known as Hamiltonian simulation (HS). HS has been invented to simulate quantum systems and is among the most promising applications of QCs with the potential for exponential speedup over classical computers (Low and Chuang (2019)). It has recently been demonstrated in a handful of studies that classical wave problems can be mapped to a Schrödinger equation so that HS can be harnessed to simulate Maxwell’s equations (Jin et al. (2022, 2023)), homogeneous acoustic waves, and mass-spring models (Suau et al. (2021); Costa et al. (2019); Babbush et al. (2023)). This study exploits HS to solve the 1-D elastic wave equation in heterogeneous media. Additionally, we perform initial experiments on actual QCs.
More specifically, we extend numerical wave propagation via HS (Costa et al., 2019; Suau et al., 2021) to heterogeneous media, a necessary step toward FWI on QCs. For this, we employ an analytical Cholesky decomposition, which facilitates the transformation of the elastic wave equation with heterogeneous medium properties to a Schrödinger equation. We proceed with the implementation of our method on the QC IBM Brisbane, in order to solve a small-scale problem and explain current challenges. These include systematic errors caused by noise and decoherence rates, which are characteristic for current NISQ QCs. Anticipating future developments, we finally delve into the potential scalability of an optimal HS algorithm on future quantum hardware equipped with error correction. By recognizing our problem as a $d$-sparse HS (Low and Chuang (2019); Babbush et al. (2023)), we establish that the algorithm is capable of providing exponential speedup compared to computations on classical computers.
To conclude the study, we briefly elaborate on perspectives to resolve the read-out problem, arising from the probabilistic nature of quantum mechanics. We further derive a continuous version of the classical-to-quantum map and discuss how this paves the way for employing other discretization methods. Finally, we discuss the extensions and limitations of our method and how it relates to other quantum algorithms.
@@ 2 A brief summary of quantum computing
To set the stage, we begin with a condensed review of the tensor product formalism and the quantum circuit model that should be accessible to a geophysical audience. This framework is inherently dictated by the postulates of quantum mechanics and forms the fundamental programming model of quantum computing. A more general formalism of density matrices, as well as more in-depth introductions, can be found in the literature (Nielsen and Chuang (2010); Watrous (2018); Adedoyin et al. (2022); Benenti et al. (2019)).
@@ 2.1 Quantum particles and state space
Consider a single quantum particle characterized by its quantized properties, such as the atomic energy level. This particle can occupy $N$ different orthogonal basis states, which span the state space of the particle. To the basis states, we assign labels $\Sigma=\{0,\dots,N-1\}$, collectively referred to as the label set of size $|\Sigma|=N$. A particle with $N$ basis states is called an $N$-level particle. The current state of the observable is represented by the state vector |$\ket{\psi}=[\alpha_{0},\dots,\alpha_{N-1}]^{T}\in\mathbb{C}^{N}$, where $\alpha_{k}\in\mathbb{C}$ is the probability amplitude corresponding to the $k$-th basis state |$\ket{k}$, written in the bra-ket or Dirac notation that is ubiquitous in quantum mechanics. It encapsulates all information about the individual particles and their number and concisely represents exponentially large state vectors. The ket |$\ket{a}$ is a column vector, whereas the bra $\bra{b}$ | is a row vector with complex conjugate transpose $\bra{b}^{\dagger}=\ket{b}$. Bra and ket form the bra-ket $\braket{b}{a}$, which is the scalar product |$\ket{b}^{\dagger}\cdot\ket{a}$. The main difference, compared to the common vector notation in linear algebra, is that the content of the bra/ket does not represent individual vector entries but some convenient collective label. The state vector |$\ket{\psi}$ can be expressed in terms of the basis states |$\ket{k}$ with $k=0,...,N-1$ as the linear combination |$\ket{\psi}=[\alpha_{0},\dots,\alpha_{N-1}]^{T}=\alpha_{0}\ket{0}+\alpha_{0}\ket{1}+\dots+\alpha_{N-1}\ket{N-1}\in\mathbb{C}^{N}$, which is called a superposition state.
@@
where |$\ket{\psi_{j}}\in\mathbb{C}^{N_{j}}$ is the state of the $j^{\text{th}}$ particle, and $N_{j}$ is the number of states that it can occupy. We reuse the symbols $N$ and $\psi$, as a single-particle system is a special case of a multi-particle system. The symbol $\otimes$ denotes the Kronecker or tensor product. For two arbitrary states $\mathbf{u}$ and $\mathbf{v}$, it is defined as
@@
Any quantum state must be normalized, i.e., satisfy the normalization condition
@@
Combining normalized states with the $\otimes$ product, the resulting composite state is normalized as well. If $\alpha_{k}\neq 0$ for at least two different indices $k$, then the state |$\ket{\psi}$ simultaneously holds multiple potential outcomes $\Omega=\{\ket{k}:\alpha_{k}\neq 0\}$ and is called a superposition state. In contrast to classical systems, a measurement of the quantum system causes the state |$\ket{\psi}$ to collapse from the superposition into one of the possible basis states |$\ket{k}\in\Omega$ randomly with probability $P(k)=|\alpha_{k}|^{2}$, which explains the term probability amplitude used for the coefficients $\alpha_{k}$. The type of measurement just introduced is a full-system measurement in the standard basis, which is the simplest case; we do not deal with more general cases, which can be found in the literature at the top of this section. Measurements allow us to read the label of the collapsed state in the form of classical bits, forming the only interface between quantum and classical information.
@@ 2.3 Hamiltonian simulation
with the initial state |$\ket{\psi(0)}$. The time evolution operator $\mathbf{U}(t)=\exp(-i\mathbf{H}t)$ is unitary. Our approach to elastic wave propagation on a QC rests on the quantum or Hamiltonian simulation (HS) approach (e.g., Georgescu et al., 2014; Nielsen and Chuang, 2010), which amounts to the direct evaluation of (9). HS was originally introduced to model multi-particle quantum dynamic systems using QCs. Due to the aforementioned exponential space growth, quantum dynamics modeling becomes prohibitively complex on classical computers. In contrast, by being quantum systems themselves, QCs naturally feature the same state space expansion, potentially making multi-particle simulations possible.
For initialization, HS requires the application of a unitary operation $\mathbf{V}$ that sets the desired initial state, i.e. |$\ket{\psi(0)}=\mathbf{V}\ket{00..00}$. Subsequently, the unitary time evolution operator $\mathbf{U}(t)$, constructed using basic gates, is applied to |$\ket{\psi(0)}$,
@@
The main challenge of HS is that the matrix exponential $\mathbf{U}(t)=\exp(-i\mathbf{H}t)$ is notoriously difficult to evaluate exactly. Hence, the application of $\mathbf{U}(t)$ is replaced by some unitary $\mathbf{U}_{\textrm{A}}(t)$ that evolves the initial state to the final one with the desired error $\varepsilon$, $||\mathbf{U}(t)\ket{\psi(0)}-\mathbf{U}_{\textrm{A}}(t)\ket{\psi(0)}||<\varepsilon$, and can be implemented efficiently using basic quantum gates, with circuit depth growing along with $1/\varepsilon$.
@@
Other approaches can be found in the literature, with efficiency varying based on properties of the given $\mathbf{H}$. These include methods based on the Trotter-Suzuki Decomposition (e.g., Trotter, 1959; Suzuki, 1976, 1991; Hatano and Suzuki, 2005; Dhand and Sanders, 2014; Yi and Crosson, 2022) and the recent qubitization approach (Low and Chuang, 2019), discussed more in detail in appendix B.
It has been shown that HS can achieve up to exponential speedup compared to analogous classical algorithms (e.g., Berry et al., 2007; Childs et al., 2018; Daley et al., 2022; Babbush et al., 2023; Low and Chuang, 2019). We will discuss this matter as well as the question of utilizing the final state in section 5.
@@ 3 Elastic waveform propagation using HS
with time $t$, the wave field $u(x,t)$, density $\rho(x)$ and elastic modulus $\mu(x)$. Furthermore, the wave field is subject to the initial conditions $u_{0}=u(x,0)$ and $v_{0}=\partial_{t}u(x,0)$. We introduce a first-order finite-difference discretization in space with $2^{n-1}$ grid points, each with position and velocity, and a grid spacing of $\Delta x=W/(2^{n-1}-1)$, thereby approximating the continuous wave field $u(x,t)$ by the discrete wave field vector $\mathbf{u}(t)=[u(x_{0},t),\dots,u(x_{2^{n-1}-1},t)]^{T}$. In section 5.3, we discuss the potential use of other discretization methods. We further define the positive definite mass matrix $\mathbf{M}$ that serves as a discrete representation of the density distribution,
@@
This implementation naturally results in a Dirichlet and a Neumann boundary condition on the right and left, respectively, assuming two ghost nodes. In the interest of a smooth derivation, we defer the case of arbitrary boundary conditions to appendix C. With these definitions, the space-discretized wave equation reads
@@
To solve (16) on a QC, it needs to be transformed into the Schrödinger equation (8). For this, we first apply a transformation that utilizes the positive definite, diagonal square root matrix $\mathbf{M}^{1/2}$, which produces the mass-transformed wave field $\tilde{\mathbf{u}}=\mathbf{M}^{1/2}\mathbf{u}$ and the mass-transformed stiffness matrix $\tilde{\mathbf{K}}=\mathbf{M}^{-1/2}\mathbf{K}\mathbf{M}^{-1/2}$. The latter inherits the property of being symmetric negative-definite from $\mathbf{K}$. We proceed by transforming the second-order system into a first-order system
@@
where $\tilde{\mathbf{v}}=\partial_{t}\tilde{\mathbf{u}}=\mathbf{M}^{1/2}\partial_{t}\mathbf{u}\in\mathbb{R}^{2^{n-1}}$ is the mass-transformed velocity vector and $\mathbf{Q}$ the impedance matrix. While (17) is already closer to the Schrödinger equation than (16), it remains to symmetrize $\mathbf{Q}$ and obtain the imaginary unit $i$ on the left-hand side. For this, we multiply (17) by an arbitrary invertible transformation $\mathbf{T}\in\mathbb{R}^{2^{n}\times 2^{n}}$, which results in
@@
which are typically sparse in seismic applications thanks to the spatial localization of sources. Furthermore, owing to the localized interactions of the discretized degrees of freedom, both the mass and stiffness matrices are sparse, which implies that $\mathbf{Q}$ is sparse, too. Eq. (19) can, in principle, be solved by a QC, and it is isomorphic to the discretized elastic wave equation (16) because $\mathbf{T}$ is invertible. The actual elastic wave field at all times can, therefore, be reconstructed via